# Generalized Moving Least-Squares for Solving Vector-valued PDEs on Unknown Manifolds

### v1

@@ 5.2 Poisson problems
Figure 3: Poisson problems on 2D torus in $\mathbb{R}^{3}$. Random data with 6 trials. $K=51$ nearest neighbors are used. In panels (a) and (b), shown are the average FEs of the approximate Bochner Laplacian $\mathbf{L}_{B}$ using extrinsic and intrinsic methods, respectively. In panels (c) and (d), shown are the average IEs of the solutions of the Poisson problem (36).
In this subsection, we consider solving the following vector-valued Poisson problem for $u\in\mathfrak{X}(M)$,
$(a-\Delta_{B})u=f,\quad\mathbf{x}\in M,$ (36)
where $a=1>0$ and $f\in\mathfrak{X}(M)$ are defined such that the problems are well-posed. Here, we notice that the vector fields $f$ and $u$ must live in the tangent bundle $TM$ of the smooth manifold $M$. For the setup of an unknown manifold, $f$ is replaced by $\mathbf{\hat{T}}^{\top}F$ in our numerical implementation where $\mathbf{\hat{T}}=[\bm{\hat{t}}_{1},\bm{\hat{t}}_{2}]$ corresponds to the approximate tangent vectors and $F\in\mathfrak{X}(\mathbb{R}^{3})$ is a vector field in ambient space. Thus, the vector field $\mathbf{\hat{T}}^{\top}F$ and the approximate solution $\hat{u}=(a-\mathbf{L}_{B})^{-1}(\mathbf{\hat{T}}^{\top}F)$ both live in the approximate tangent bundle $\widehat{TM}:=\cup_{\mathbf{x}\in M}\widehat{T_{\mathbf{x}}M}=\cup_{\mathbf{x}\in M}\mathrm{Span}\{\bm{\hat{t}}_{1}(\mathbf{x}),\bm{\hat{t}}_{2}(\mathbf{x})\}$.
@@
In this numerical experiment, we consider solving the Poisson problem (36) on a 2D torus,
@@
$\mathbf{x}=\left[\begin{matrix}(R+r\cos\theta)\cos\phi\\ (R+r\cos\theta)\sin\phi\\ r\sin\theta\end{matrix}\right],\quad 0\leq\theta<2\pi,\quad 0\leq\phi<2\pi,$ (37)
where $R$ is the distance from the center of the tube to the center of the torus and $r<R$ is the distance from the center of the tube to the surface of the tube. In our experiment, we set $R=2$ and $r=1$. The true solution $u$ is set to be $u=\sin\theta\sin\phi\frac{\partial}{\partial\theta}+\sin\theta\cos\phi\frac{\partial}{\partial\phi}$ $\in\mathfrak{X}(M)$. In our computation, the manufactured $f$ is calculated by $f:=\mathbf{\hat{T}}^{\top}F=\mathbf{\hat{T}}^{\top}(a-\bar{\Delta}_{B})U$, where $(a-\bar{\Delta}_{B})U$ is an extension of $(a-\Delta_{B})u$ as defined in (26).
Numerically, the points $\{\mathbf{x}_{i}\}_{i=1}^{N}$ are generated from the parametrization (37) using randomly sampled $\{\theta_{i},\phi_{i}\}_{i=1}^{N}$ with the uniform distribution on $\left[0,2\pi\right)\times\left[0,2\pi\right)$. We fix $K=51$ nearest neighbors in each local stencil. To verify the convergence over a number of points $N=[1600,3200,6400,12800,25600]$, we define the forward error (FE) and the inverse error (IE) in the ambient space $\mathbb{R}^{3}$, respectively, as
$\textbf{FE}=\max_{1\leq i\leq N}\|\hat{\textbf{T}}(\mathbf{x}_{i})(\mathbf{L}_{B}\mathbf{u})_{i}-\textbf{T}(\mathbf{x}_{i})\Delta_{B}u(\mathbf{x}_{i})\|_{2},\mathrm{\ \ \ }\textbf{IE}=\max_{1\leq i\leq N}\|(\hat{\textbf{T}}\hat{u})(\mathbf{x}_{i})-(\mathbf{T}u)(\mathbf{x}_{i})\|_{2}$
where $\|\cdot\|_{2}$ is the Euclidean distance, the true solution $\mathbf{u}=(u(\mathbf{x}_{1}),\ldots,u(\mathbf{x}_{N}))^{\top}\in\mathbb{R}^{2N\times 1}$ is a column vector, $(\mathbf{L}_{B}\mathbf{u})_{i}\in\mathbb{R}^{2\times 1}$ corresponds to the approximate vector value of $\Delta_{B}u$ at $\mathbf{x}_{i}$, $\Delta_{B}u(\mathbf{x}_{i})\in\mathbb{R}^{2\times 1}$ is the representation in the global basis at $\mathbf{x}_{i}$, $\hat{u}$ is the approximate solution, $\mathbf{T}(\mathbf{x})$ $=[\bm{t}_{1}(\mathbf{x}),$ $\bm{t}_{2}(\mathbf{x})]$ $\in\mathbb{R}^{3\times 2}$ and $\mathbf{\hat{T}}(\mathbf{x})=[\bm{\hat{t}}_{1}(\mathbf{x}),\bm{\hat{t}}_{2}(\mathbf{x})]\in\mathbb{R}^{3\times 2}$ correspond to the true and the approximate global bases, respectively. Here we push both tangent vectors on (approximate) manifolds forward to tangent vectors in $\mathbb{R}^{3}$ to well-define the FE and IE in the Euclidean norm. Notice that if the manifold is known with given global bases $\mathbf{T}(\mathbf{x})$, then $\mathbf{T}(\mathbf{x})$ can be ignored in the definition of FE and IE since $\mathbf{T}:T_{\iota^{-1}\mathbf{x}}O\rightarrow T_{\mathbf{x}}M$ is an isometric mapping where $\iota$ is the local parametrization as defined in Section 4.1.
@@
In Fig. 3 , we plot the averages of $\mathbf{FE}$s and $\mathbf{IE}$s over 6 independent trials as functions of $N$. One can see that the $\mathbf{FE}$s using degree $l=2,3,4,5$ decrease on the order of $N^{-(l-1)/2}$ for both the extrinsic and intrinsic methods. This numerical error rate is consistent with the GMLS error bound in (7) since the Bochner Laplacian is involved with the second-order derivative of component functions. One can further observe that the $\mathbf{IE}$s for $l=3$ and $l=5$ decay with respective rates $N^{-1}$ and $N^{-2}$ as the GMLS theory predicts. Moreover, the $\mathbf{IE}$s possess the super-convergence for degree $l=2$ and $l=4$, that is, the $\mathbf{IE}$s decay on the order of $N^{-1}$ and $N^{-2}$, respectively, which are half-order faster than the corresponding $\mathbf{FE}$s. This super-convergence phenomenon has also been observed in the scalar case for the Laplace-Beltrami operator of functions [33, 26].
@@ 5.3 Linear time-dependent equations
$u_{t}=\nu\Delta_{B}u+f,\quad(\mathbf{x},t)\in M\times\left(0,T\right]\text{ with }u(\cdot,0)=u_{0}(\cdot),$ (38)
@@
where $-\pi/2\leq\theta\leq\pi/2$, $-\pi\leq\phi<\pi$, $r=3.91/3.39$, $c_{0}=0.81/3.39$, $c_{2}=7.83/3.39$ and $c_{4}=-4.39/3.39$, which is the same as the RBC in [17]. The second one is a perturbed sphere (PSP) with the parametrization:
@@
where $0\leq\theta\leq\pi$, $0\leq\phi<2\pi$ and $r(\theta,\phi)=1+0.05(1-\cos 4\theta)\sin 2\phi$.
In our numerical experiments, we set the external forcing $f(\textbf{x},t):=\hat{\textbf{T}}^{\top}\textbf{x}$ in both examples to be time-independent. The initial condition is $u_{0}(\textbf{x}):=0.2\hat{\textbf{T}}^{\top}\textbf{x}$ for RBC and $u_{0}(\textbf{x}):=0.16\hat{\textbf{T}}^{\top}\textbf{x}$ for PSP. Here, the coefficients in $u_{0}$ are specified such that the vector-field solutions on a scale of $0.1$.
(a) RBC (extr.) (b) RBC (intr.) (c) PSP (extr.) (d) PSP (intr.)
@@
Figure 4: Linear vector diffusion equation on RBC and PSP. Time step $\Delta t=0.001$ and $K$-nearest neighbor $K=61$. Shown in the first two rows are the randomly sampled scatter data points on RBC (the first row) and PSP (the second row) with the color bar representing the magnitude of the vector fields $\hat{u}$ and the red arrows indicating the direction of the vector fields $\hat{u}$. In the third row, we show the convergence examination of the average SEs at $t=0.05$ with error bars obtained from the standard deviations over 6 independent trials.
To solve the time evolution equation (38), we use the GMLS method for the space discretization and the second-order explicit Runge-Kutta method for the time discretization. For both examples, we set $K=61$ in $K$ nearest neighbors and the time step $\Delta t=0.001$. The data are randomly sampled from the intrinsic coordinate and then mapped onto the surface using the above parametrization. To examine the convergence, we define the time-dependent solution error (SE) between true solution $u$ and approximate solution $\hat{u}$:
$\textbf{SE}=\max_{1\leq i\leq N}\|\hat{\textbf{T}}(\mathbf{x}_{i})\hat{u}(\mathbf{x}_{i},t)-\mathbf{T}(\mathbf{x}_{i})u(\mathbf{x}_{i},t)\|_{2}.$
In Fig. 4, we show our numerical results on RBC and PSP. In the first two rows, we show the time evolution of the extrinsic GMLS solutions on surfaces using $N=20000$ data points at time $t=[0,1,4,10]$. After a sufficiently long time $t=10$, we observe that the solution has arrived at its steady state given by the solution of the corresponding Poisson equation. In the last row, we show the average SEs for the numerical solutions as functions of $N$ at time $t=0.05$. To compute the reference solution, we apply the same discretization schemes over $N=80000$ randomly sampled points, with polynomial degree $l=5$, and analytic true tangent space spanned by $\bm{t}_{1},\bm{t}_{2}$. To verify the convergence, we compute the numerical solutions based on 6 independent subsampling from the $80000$ randomly sampled data points with $N=[7000,14000,28000,56000]$ on RBC and $N=[5000,10000,20000,40000]$ on PSP. From the last row of Fig. 4, we can still see the super-convergence phenomena for both extrinsic formulation and intrinsic formulation in the two examples. We have also examined such super-convergence phenomena using a second-order backward differentiation formula (BDF2) [9] for time-discretization.
In addition, we also applied our GMLS approach to bumpy sphere examples which have the same parametrization form as in (40) but with the radial function $r(\theta,\phi)=1+0.05(1-\cos w_{\theta}\theta)\sin w_{\phi}\phi$ having perturbations from higher wavenumbers $w_{\theta},w_{\phi}$. However, we found that the convergence rate becomes slower than that predicted by the GMLS theory when $w_{\theta}=6,w_{\phi}=6$ and the Laplacian matrix even suffers from the eigenvalue instability when $w_{\theta},w_{\phi}$ become even larger [not shown here]. Notice that larger wavenumbers $w_{\theta}$ and $w_{\phi}$ give us a bumpy sphere with larger local curvatures. From the above numerical results, we believe that the current GMLS approach with the specified weight function in (3) can stabilize the vector Laplacians when manifolds are smooth with small curvatures. Nevertheless, when curvatures become large, the current GMLS approach may not provide a stable approximation to vector Laplacians.
@@ 5.4 Viscous Burgers’ equation
$u_{t}+\nabla_{u}u=\nu\Delta_{B}u+f,\quad(\mathbf{x},t)\in M\times\left(0,T\right]\text{ with }u(\cdot,0)=u_{0}(\cdot),$ (41)
where $u(\cdot,t)\in\mathfrak{X}(M)$ represents a velocity field and $\nu=0.01$ is a viscosity parameter. In this example, the parametrization of the torus is given in (37) and the points are sampled in the same way as those in subsection 5.2. For the time discretization, we use the second-order Runge-Kutta method with $\Delta t=0.001$.
@@
Figure 5: (Color online) Linear vector diffusion equation and nonlinear Burgers’ equation on 2D torus. Time step $\Delta t=0.001$ and $K$-nearest neighbor $K=71$. In the first two rows, we show the time evolution of the extrinsic GMLS solutions of linear (the first row) and nonlinear (the second row) equations using $N=6400$ randomly sampled data points on the torus, with the color bar representing the magnitude of the vector fields $\hat{u}$ and the red arrows indicating the direction of the vector fields $\hat{u}$. Shown in the third row is the convergence examination of the average SEs with error bars obtained from the standard deviations over 6 independent trials at $t=0.05$.
In the first two rows of Fig. 5, we show the comparison of evolution of the extrinsic GMLS solutions between the linear equation (without covariant derivative in (41)) and the nonlinear viscous Burgers’ equation (with covariant derivative). For both equations, we set the same initial condition $u_{0}={0}$ to be a zero vector field and use the same external forcing $f:=-\nu\Delta_{B}w$ with $w=0.05\sin\theta\sin\phi\frac{\partial}{\partial\theta}+0.05\sin\theta\cos\phi\frac{\partial}{\partial\phi}$. In our numerical implementation for the unknown manifold, we project the forcing onto the approximate tangent bundle by multiplying by $\mathbf{\hat{T}}^{\top}$ as the same in subsection 5.2. One can see from the first two rows of Fig. 5 that the solutions behave similarly for a short time ($t=50$) while they behave differently at their steady states for sufficiently long time ($t=100$). In the third row of Fig. 5, we verify the convergence of the proposed methods for solving the linear equation and the nonlinear Burgers’ equation. The true solution is set to be $u=0.05\cos t\sin\theta\sin\phi\frac{\partial}{\partial\theta}$ $+0.05\cos t\sin\theta\cos\phi\frac{\partial}{\partial\phi}$ for both equations. One can see that the average SEs at time $t=0.05$ over 6 independent trials decrease nearly on the order of $N^{-(l-1)/2}$ for degree $l=2,3,4,5$ as the GMLS theory predicts in (7). Here, we notice that the super-convergence phenomenon, observed previously in Section 5.3 (where the viscosity parameter $\nu=0.1$ was set to be relatively large), is not obviously observed here for relatively small viscosity $\nu=0.01$ in both the linear and nonlinear regimes.
In addition, we numerically simulated the regime for the external forcing with large amplitude. We observed that shock waves are first formed in the solution and then the solution suffers from numerical instability [not shown here]. We also numerically observed such instability results by using the Crank-Nicolson/Adams-Bashforth scheme [29]. The reason is that in the local hypothesis space only containing multi-variate polynomials, it cannot capture the discontinuous solution such as shock waves well.

### v2

@@ 5.2.2 Bochner Laplacian on 2D torus in $\mathbb{R}^{9}$
Figure 6: Screened Poisson problems on 2D torus in $\mathbb{R}^{9}$. Random data with 4 trials. $K=55$ nearest neighbors are used. In panels (a) and (b), shown are the average FEs of the approximate Bochner Laplacian $\mathbf{L}_{B}$ using extrinsic and intrinsic methods, respectively. In panels (c) and (d), shown are the average IEs of the solutions of the screened Poisson problem (49).
@@
In this numerical experiment, we consider solving the screened Poisson problem (49) on a 2D torus in $\mathbb{R}^{9}$,
@@
$\mathbf{x}=\left[\begin{matrix}(R+\cos\theta)\cos\phi\\ (R+\cos\theta)\sin\phi\\ \vdots\\ (1/4)(R+\cos\theta)\cos 4\phi\\ (1/4)(R+\cos\theta)\sin 4\phi\\ \sqrt{\sum_{i=1}^{4}(1/i)^{2}}\sin\theta\end{matrix}\right]\in\mathbb{R}^{9},\quad 0\leq\theta<2\pi,\quad 0\leq\phi<2\pi,$ (50)
where $R=2$ is the distance from the center of the tube to the center of the torus. The true solution $\bm{u}$ is set to be $\bm{u}=\sin\theta\sin\phi\frac{\partial}{\partial\theta}+\sin\theta\cos\phi\frac{\partial}{\partial\phi}$ $\in\mathfrak{X}(M)$. The point cloud $\{\mathbf{x}_{i}\}_{i=1}^{N}$ is generated using the parametrization (50), where $\{\theta_{i},\phi_{i}\}_{i=1}^{N}$ are randomly sampled from a uniform distribution on $\left[0,2\pi\right)\times\left[0,2\pi\right)$. We fix $K=50$ in $K$-nearest neighbors.
@@
In Fig. 6 , we plot the averages of $\mathbf{FE}$s and $\mathbf{IE}$s over 4 independent trials as functions of $N$. One can see that the $\mathbf{FE}$s using degree $l=2,3,4,5$ decrease on the order of $N^{-(l-1)/2}$ for both the extrinsic and intrinsic methods. This numerical error rate is consistent with the GMLS error bound in (7) since the Bochner Laplacian is involved with the second-order derivative of component functions. One can further observe that the $\mathbf{IE}$s for $l=3$ and $l=5$ decay with respective rates $N^{-1}$ and $N^{-2}$ as the GMLS theory predicts. Moreover, the $\mathbf{IE}$s possess the super-convergence when degree $l=2$ and $l=4$, that is, the $\mathbf{IE}$s decay on the order of $N^{-1}$ and $N^{-2}$, respectively, which are half-order faster than the corresponding $\mathbf{FE}$s. This super-convergence phenomenon has also been observed in the scalar case for the Laplace-Beltrami operator of functions [37, 28].
5.2.3 3D flat torus in $\mathbb{R}^{12}$
We consider a $3$-dimensional manifold embedded in $\mathbb{R}^{12}$ with the following parameterization,
$\mathbf{x}=\frac{1}{\sqrt{1^{2}+2^{2}}}\left(\begin{array}[]{cccc}\cos(\phi_{1}),&\sin(\phi_{1}),&\cos(2\phi_{1}),&\sin(2\phi_{1}),\\ \cos(\phi_{2}),&\sin(\phi_{2}),&\cos(2\phi_{2}),&\sin(2\phi_{2}),\\ \cos(\phi_{3}),&\sin(\phi_{3}),&\cos(2\phi_{3}),&\sin(2\phi_{3})\end{array}\right),$ (51)
with $0\leq\phi_{1},\phi_{2},\phi_{3}<2\pi$. The Riemannian metric is given by a $3\times 3$ identity matrix $\mathbf{I}_{3}$. The true solution $\bm{u}$ is set to be $\bm{u}=\sin(\phi_{1})\sin(\phi_{2})\frac{\partial}{\partial\phi_{1}}+\sin(\phi_{2})\sin(\phi_{3})\frac{\partial}{\partial\phi_{2}}+\sin(\phi_{3})\cos(\phi_{1})\frac{\partial}{\partial\phi_{3}}$. Numerically, the points $\{\mathbf{x}_{i}\}_{i=1}^{N}$ are generated from the parametrization (51) using randomly sampled $\{\phi_{1}(\mathbf{x}_{i}),\phi_{2}(\mathbf{x}_{i}),\phi_{3}(\mathbf{x}_{i})\}_{i=1}^{N}$ with the uniform distribution on $\left[0,2\pi\right)^{3}$. We use the degree $l=3,4$ and $K=75$ nearest neighbors for illustrating the convergence of solutions.
In Fig. 7, we illustrate the convergence of $\textbf{IE}s$ for both the extrinsic and intrinsic GMLS methods. From Figs. 7(a) and (b), we observe that both methods provide a stable and convergent approximation to the solution for both known and unknown manifold setups. Additionally, the extrinsic method exhibits nearly the same errors between known and unknown manifold setups. However, for the intrinsic method, the error for the unknown manifold is much larger than that for the known manifold while the convergence rates hold. Upon further examination, we identify that the large error gap in the intrinsic method for this 3D example stems from the local regression approximation of Monge parametrization in Step 8 of Algorithm 1. Specifically, the tangent space of the local GMLS approximation of Monge parametrization ($\mathrm{span}\{\partial_{1},\partial_{2},\partial_{3}\}$) deviates from that of the manifold ($\mathrm{span}\{\bm{t}_{1},\bm{t}_{2},\bm{t}_{3}\}$) at some neighboring points in the local stencil $S_{\mathbf{x}_{0}}$. To improve the result for the intrinsic method, one potential approach is to use an interpolation (see e.g. [15, 29]) instead of a regression for approximating the local manifold. This suggests a direction for future research.
(a) (extr.) IE (b) (intr.) IE (c) Pointwise Absolute Error
Refer to caption Refer to caption Refer to caption
Figure 7: Screened Poisson problems on 3D flat torus in $\mathbb{R}^{12}$. $K=75$ nearest neighbors and degree $l=3,4$ are used. In panels (a) and (b), shown are the IEs of the solutions using our extrinsic and intrinsic methods, respectively. In panel (c), shown is the pointwise absolute error of the solution using extrinsic method for an unknown manifold setup with $N=51200$.
@@ 5.3 Linear time-dependent equations
$\bm{u}_{t}=\nu\Delta_{B}\bm{u}+\bm{f},\quad(\mathbf{x},t)\in M\times\left(0,T\right]\text{ with }\bm{u}(\cdot,0)=\bm{u}_{0}(\cdot),$ (52)
@@
where $-\pi/2\leq\theta\leq\pi/2$, $-\pi\leq\phi<\pi$, $r=3.91/3.39$, $c_{0}=0.81/3.39$, $c_{2}=7.83/3.39$ and $c_{4}=-4.39/3.39$, which is the same as the RBC in [18]. The second example is a bumpy sphere (BSP) with the parametrization:
@@
where $0\leq\theta\leq\pi$, $0\leq\phi<2\pi$ and $r(\theta,\phi)=1+0.2(\sin^{7}3\theta)(\sin 4\phi)$.
In our numerical experiments, we set the external forcing $\bm{f}(\textbf{x},t):=\hat{\textbf{T}}^{\top}\textbf{x}$ in both examples to be time-independent. The initial condition is $\bm{u}_{0}(\textbf{x}):=0.2\hat{\textbf{T}}^{\top}\textbf{x}$ for RBC and $\bm{u}_{0}(\textbf{x}):=0.1\hat{\textbf{T}}^{\top}\textbf{x}$ for BSP. To solve the time evolution equation (52), we use the GMLS method for the space discretization and the second-order explicit Runge-Kutta method for the time discretization. The data are randomly sampled from the intrinsic coordinate and then mapped onto the surface using the above parametrization. To examine the convergence, we define the time-dependent solution error (SE) between true solution $\bm{u}$ and approximate solution $\hat{\mathbf{u}}$:
$\textbf{SE}=\max_{1\leq i\leq N}\|\hat{\textbf{T}}(\mathbf{x}_{i})\hat{\mathbf{u}}(\mathbf{x}_{i},t)-\mathbf{T}(\mathbf{x}_{i})\bm{u}(\mathbf{x}_{i},t)\|_{2}.$
(a) RBC (extr.) (b) RBC (intr.) (c) BSP (extr.) (d) BSP (intr.)
@@
Figure 8: Linear vector diffusion equation on RBC and BSP. For both examples, we set $K=60$ in $K$ nearest neighbors. The time step $\Delta t=0.001$ for RBC and $\Delta t=0.0001$ for BSP. Shown in the first two rows are time evolution of the extrinsic GMLS solutions on RBC (the first row) and BSP (the second row) with the color bar representing the magnitude of the vector fields $\hat{\mathbf{u}}$ and the red arrows indicating the direction of the vector fields $\hat{\mathbf{u}}$. We use $N=20000$ randomly sampled data points, degree $l=2$ for RBC and $N=160000$ randomly sampled data points, degree $l=4$ for BSP. In the third row, we show the convergence examination of the average SEs at $t=0.05$ with error bars obtained from the standard deviations over 6 independent trials.
In Fig. 8, we show our numerical results on RBC and BSP. The first two rows illustrate the time evolution of the extrinsic GMLS solutions on surfaces using $N=20000$ data points for RBC and $N=160000$ data points for BSP at times $t=[0,1,4,10]$. After a sufficiently long time ($t=10$), the solution reaches its steady state, corresponding to the solution of the associated Poisson equation. The last row shows the average SEs for the numerical solutions as functions of $N$ at $t=0.05$. To compute the reference solution, we apply the same discretization schemes over $N=80000$ randomly sampled points for RBC and $N=640000$ for BSP, with polynomial degree $l=5$, and analytic true tangent space spanned by $\bm{t}_{1},\bm{t}_{2}$. To verify the convergence, we compute the numerical solutions based on 6 independent subsamplings with $N=[7000,14000,28000,56000]$ on RBC and $N=[20000,40000,80000,160000,320000]$ on BSP. From the last row of Fig. 8, we can still see the super-convergence phenomena for both extrinsic formulation and intrinsic formulation in the two examples. We have also examined such super-convergence phenomena using a second-order backward differentiation formula (BDF2) [9] for time-discretization.
@@ 5.4 Viscous Burgers’ equation
$\bm{u}_{t}+\nabla_{\bm{u}}\bm{u}=\nu\Delta_{B}\bm{u}+\bm{f},\quad(\mathbf{x},t)\in M\times\left(0,T\right]\text{ with }\bm{u}(\cdot,0)=\bm{u}_{0}(\cdot),$ (55)
where $\bm{u}(\cdot,t)\in\mathfrak{X}(M)$ represents a velocity field and $\nu=0.1$ is a viscosity parameter. In this example, the parametrization of the torus is given in (48) and the points are sampled in the same way as those in subsection 5.1. For the time discretization, we use the second-order Runge-Kutta method with $\Delta t=0.0001$. We also examine the following result using Crank-Nicolson/Adams-Bashforth scheme [31].
@@
Figure 9: Linear vector diffusion equation and nonlinear Burgers’ equation on 2D torus in $\mathbb{R}^{3}$. Time step $\Delta t=0.0001$ and $K$-nearest neighbor $K=70$. In the first two rows, we show the time evolution of the extrinsic GMLS solutions of linear (the first row) and nonlinear (the second row) equations using $N=6400$, degree $l=2$, with the color bar representing the magnitude of the vector fields $\hat{\mathbf{u}}$ and the red arrows indicating the direction of $\hat{\mathbf{u}}$. Shown in the third row is the convergence examination of the average SEs with error bars obtained from the standard deviations over 6 independent trials at $t=0.05$.
In the first two rows of Fig. 9, we compare the evolution of the extrinsic GMLS solutions between the linear equation (without covariant derivative) in (52) and the nonlinear viscous Burgers’ equation (with covariant derivative) in (55). For both equations, we set the same initial condition $\bm{u}_{0}={0}$ to be a zero vector field and use the same external forcing $\bm{f}:=-\nu\Delta_{B}\bm{w}$ with $\bm{w}=0.05\sin 2\theta\sin 3\phi\frac{\partial}{\partial\theta}+0.05\sin\theta\cos\phi\frac{\partial}{\partial\phi}$. In our numerical implementation for the unknown manifold, we project the forcing onto the approximate tangent bundle by multiplying by $\mathbf{\hat{T}}^{\top}$ as the same in subsection 5.2. One can see from the first two rows of Fig. 9 that the solutions exhibit similar dynamical behavior for a relatively short time ($t=10$) since the covariant derivative term has a small contribution for a short period. However, the solutions behave quite differently at their steady states for a sufficiently long time ($t=200$).
In the third row of Fig. 9, we verify the convergence of the proposed GMLS methods for solving the linear equation and the nonlinear Burgers’ equation. The true solution is set to be $\bm{u}=0.05\cos t\sin\theta\sin\phi\frac{\partial}{\partial\theta}$ $+0.05\cos t\sin\theta\cos\phi\frac{\partial}{\partial\phi}$ for both equations. One can see that the average SEs at time $t=0.05$ over 6 independent trials decrease at least on the order of $N^{-(l-1)/2}$ when degree $l=2,3,4,5$ as the GMLS theory predicts in (7). Here, the super-convergence can still be observed using the extrinsic method as seen in previous Section 5.2 and Section 5.3 while it can be barely seen using the intrinsic method.